OpenStudy (anonymous):

Write as a number in standard notation: 3.58 x 10-3

OpenStudy (cggurumanjunath):

3.58/1000=0.00358

OpenStudy (anonymous):

thank you

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

Write as a number in standard notation: 8.713 x 104????????

OpenStudy (cggurumanjunath):

can u try ! i will help u .

OpenStudy (anonymous):

I got to go I have a another test that I have to do right now. but thank you for your help

OpenStudy (cggurumanjunath):

8.713*10^4=87130.000

OpenStudy (cggurumanjunath):

see it is very simple 10^4 so write 8.713*10^4 as 8.7130000 then shift 4 decimal places to the right we get 8.713*10^4=87130.000 is it clear @Autumn,moore 8.713*10^4=87130